OpenStudy (anonymous):

It is possible for a quadratic equation to have no real-number solutions. Solve. u ^ 2 + 2u - 4 = 0

OpenStudy (anonymous):

\[u ^ 2 + 2u - 4 = 0\]\[Δ=2^2-4\times(-4)=20\]\[x_1,_2=(-2\pm \sqrt{20})/2\times1\] \[x=-1+\sqrt{5}\]or\[x=-1-\sqrt{5}\]

OpenStudy (chaise):

It is possible for the quadratic equation to have no real solution, if b^2-4ac < 0 then you will get an imaginary solution. (This also means the graph does not cut the x-axis, and i above or below it)

OpenStudy (anonymous):

u can use quadratic formula here: u =[ -2 +- sqrt(2^2 - 4 *1 * -4) ] / 2 = -1 +- sqrt 5

OpenStudy (anonymous):

thankz jimmy !